# Learning to Adapt through Bio-Inspired Gait Strategies for Versatile Quadruped Locomotion

### v2

@@ I-C2 Actuator-structural Forces – Foot Contact Forces
Animals are observed to change gait to minimise actuator-structural forces (i.e. musculoskeletal forces) [37], which in biomechanics is measured through ground reaction force, $f_{\text{grf}}$. Similarly, $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{\tau_{\%}}$ reduce $f_{\text{grf}}$ through selecting the optimal gait for minimising $\tau_{\%}$. This supports that $\tau_{\%}$ is a suitable alternative to $f_{\text{grf}}$, which is further validated by a strong correlation between them within Figure 5(b). However, not only does $\pi_{G}^{\tau_{\%}}$ maintain a trotting gait past optimal $f_{\text{grf}}$, but also the transition itself is instantaneous. In turn, $\pi_{G}^{\tau_{\%}}$ better reflects the animal data from [37] than $\pi_{G}^{\text{uni}}$. This could be explained by the metric being misinterpreted in [37]; with a high correlation between $\tau_{\%}$ and $f_{\text{grf}}$ with stability metric $c_{\text{avg}}^{\text{err}}$, it suggests instability causes transition, which is rare on flat terrain, as discussed in Section II.
@@ I-C3 Mechanical Work – External Work
Animals transition to preserve external mechanical work, $W_{\text{ext}}$, but do so before $\lambda^{W_{\text{ext}}}$, indicating relaxed minimisation (Figure 5(a)). $\pi_{G}^{\text{uni}}$ reflects this behaviour; transition occurs before $\lambda^{W_{\text{ext}}}$ yet minimal $W_{\text{ext}}$ is preserved. While $\pi_{G}^{W_{\text{ext}}}$ can reduce $W_{\text{ext}}$, its transition phase is extended over a larger range of velocities compared to animals, occurring just after $\lambda^{W_{\text{ext}}}$. In turn, this suggests that switching gaits between trotting and running offers minimal reductions in $W_{\text{ext}}$ resulting in a less definitive transition. However, $W_{\text{ext}}$ also seems to capture stability due to its high correlation with the stability metrics on $h_{\text{terr}}^{3}$ in Figure 5(b), providing insight into its reduced role in Figure 5(a) where only flat terrain is present.
@@ I-C4 Stability – Stride Duration Coefficient of Variation
Within [35] animals are shown to reduce their stride duration coefficient of variation (CV) to preserve stability as high gait periodicity indicates stability. This behaviour is presented in Figure 5(a), where animals are seen to initiate a transition phase when there is a significant increase in stride CV, consequently leading to a decrease in CV and an increase in stability. $\pi_{G}^{\text{uni}}$ inherits the same strategy as only when a spike in stride CV is experienced does a transition phase begin that results in improved stability through lowering CV. However, this is not the case with $\pi_{G}^{c^{\text{err}}}$ as it acts to reduce stride CV much more aggressively by mixing both slow, fast and auxiliary gaits which results in no clear transition phase being produced.
Only $\pi_{G}^{\text{uni}}$ consistently reflects all animal data sets and demonstrates the successful instillation of animal gait transition strategies. This also supports the notion that no singular metric can characterise animal gait selection and through unification can similar behaviour in robots arise; the minimisation of the metrics is expected and is seen across all $\pi_{G}$ policies, but the intricacies of animal gait transition strategies are only seen in $\pi_{G}^{\text{uni}}$. Additionally, we have verified this behaviour transfers to real-world deployment within Supplementary Section 2; $\pi_{G}^{\text{uni}}$ achieves an $18\%$ and $30\%$ reduction in CoT compared to trotting and running on grassy terrain while also preventing the failure cases trotting exhibits.
Figure 6: Framework deployment in real-world environments to evaluate adaptability (a) The deployment of $\pi_{L}^{\text{bio}}$ on uneven grassy terrain with manual gait selection to cycle through all gaits with a $v_{\text{x}}^{\text{cmd}}$ of $0.5$ m/s where red shaded regions indicate a transition period. For subfigures (b), (c) and (d) our framework is deployed on loose timber, muddy grass and a low-friction board, with maximum velocity magnitudes of $1$ m/s, $2$ m/s and $1.7$ m/s respectively, where the event of critical loss of stability is indicated by the red dashed line, the bottom subplot uses the same gait colour code as Figure 4, and the heatmaps show the magnitudes of each metric increasing from purple to yellow. (e) Snapshots showing how animals also use a mixture of auxiliary gaits to overcome challenging terrains, where red circles are swing feet and green circles are stance feet.
@@ I-D Adaption to Real World Terrain
Although we have instilled animal gait transition strategies, gait procedural memory, and adaptive motion into our framework, its real-world proficiency remains uncertain. Grassy terrain may trap swing feet, and the ground often features irregularities. Despite $\pi_{L}^{\text{bio}}$ only observing flat terrain during training, it can deploy all seven gaits on this terrain, as illustrated in Figure 6(a), demonstrating that gait procedural memory and adaptive motion adjustment successfully transfer to real-world environments, providing a high level of adaptability, as shown in Supplementary Video S4.
Terrain that causes states of instability presents a significant risk to the robot, hence we test the limits of our framework through deployment on loose timber, muddy grass, and a low-friction board, as presented in Figure 6(b), Figure 6(c) and Figure 6(d) respectively and consolidated in Supplementary Video S5, with an additional experiment for external perturbations included within Supplementary Section 3. Each of the presented experiments showcases an off-nominal stability recovery event; however in the nominal scenario, the framework can maintain stability without changing gaits. In the case of loose timber, critical instability is caused when one back foot slips on a plank, causing it to collide with another. In response, $\pi_{G}^{\text{uni}}$ utilises auxiliary gaits pronk and bound to recover, as depicted in Figure 6(b). This strategy of utilising the auxiliary gaits for stability recovery is seen across all experiments and reflected in animals as highlighted in Figure 6(e) where a horse is observed to utilise bounding and limping gaits to traverse down complex rock formations.
In all experiments presented, a considerable increase in a combination of $W_{\text{ext}}$, $\tau_{\%}$ and $c_{\text{avg}}^{\text{err}}$ is experienced before a gait transition, while a weaker correlation is observed with CoT; spikes in $W_{\text{ext}}$, $\tau_{\%}$ and $c_{\text{avg}}^{\text{err}}$ directly coincide or even preempt gait changes while CoT peaks lag. This is expected with $c_{\text{avg}}^{\text{err}}$ and $W_{\text{ext}}$ as they capture periodicity and base height respectively. However, this is less expected for $\tau_{\%}$ to correlate with stability; this was never a factor investigated within [37], however, the correlation observed in these experiments and in Figure 5(b) provides strong evidence that this is the case.
@@ II Discussion
From taking inspiration from animal locomotion proficiency attributes, we have developed a locomotion framework capable of traversing complex and high-risk terrain despite the robot not utilising exteroceptive sensors nor experiencing any rough terrain during the training of $\pi_{L}^{\text{bio}}$. For $\pi_{L}^{\text{bio}}$, this is achieved through including the BGS output $\bm{\beta}_{L}$ (which encodes state dependent pseudo gait procedural memory and adaptive motion adjustments) within the observation space $\bm{o}_{L}$. This proves to be effective at overcoming rough terrain within Figure 3 as without the presence of $\bm{\beta}_{L}$ within $\bm{o}_{L}$, increased failure and instability are observed. This is the equivalent of removing an animal’s cerebellum functionality, which would result in reduced limb coordination and stability [45], in turn supporting the claim that $\bm{\beta}_{L}$ effectively encodes pseudo gait procedural memory.
$\pi_{G}^{\text{uni}}$ for optimal gait selection greatly expands adaptability through instilling gait selection strategies used by animals. As demonstrated in Figure 6, $\pi_{G}^{\text{uni}}$ can maintain stability in the event of the terrain undergoing radical structural or friction coefficient adjustments. These scenarios pose risks to robots with vision systems, as they typically cannot detect ground friction or terrain changes beyond their front legs. Through the use of $\pi_{G}^{\text{uni}}$, this limitation is mitigated without implementing resource-heavy exteroceptive sensors. Comparing Figure 6(b) through 6(d) to Figure 6(e) showcases that animals and $\pi_{G}^{\text{uni}}$ utilise multiple auxiliary gaits to prevent failure. This behavior, untargeted during training, suggests that unifying these metrics encodes the intricacies of animal gait transitions.
One provoking observation is that actuator-structural forces appear to characterise instability. Considering that [37] validates by applying increased loads that could cause instability, it suggests this metric was initially misunderstood. Another observation is that despite the employed biomechanics metrics only being tested on animals completing a linear forward trajectory on flat terrain, $\pi_{G}^{\text{uni}}$ upholds animal gait transition strategies across a wide range of terrains and base velocity commands. This supports that these metrics successfully characterise gait transitions and the notion that robots can test biomechanics hypotheses, avoiding the resource, compatibility, and ethical challenges of animal testing. Moving forward, we aim to integrate our work with those that focus on Froude-free locomotion, such as [11, 12, 29] to achieve exemplary adaptability and efficiency at both Froude-characterised and Froude-free locomotion levels.
@@ III-A Control Framework Overview
At the core of this work, the Unitree A1 quadruped robot used within all experiments features $12$ degrees of freedom, $n$, which are all modeled as revolute joints, with their angular positions denoted as $\bm{q}\in{\mathbb{R}}^{n}$ and its base orientation represented as a rotation matrix $\bm{R}_{B}\in SO(\text{3})$. As discussed in Section I-A and outlined in Figure 1, both $\pi_{L}$ and $\pi_{G}$ are integrated within a control framework and supported by the SE and BGS for generation of the robot’s state data and gait references respectively. The final output of $\pi_{L}$ is target joint positions, $\bm{q}^{*}$, which then get converted into joint torques, $\bm{\tau}^{*}$, through the following PD controller to get sent to the motors,
@@
where $K_{p}$ and $K_{p}$ are the PD controller gains. Throughout this work, a constant $K_{p}=25$ N/m and $K_{d}=1$ Ns/m are used while running at $1000$ Hz, while $\pi_{L}$ and $\pi_{G}$ are run at $500$ Hz and $100$ Hz respectively.
@@ III-B Bio-inspired Gait Scheduler
The BGS primary output, $\bm{\beta}_{L}=[\bm{c}^{\text{ref}},\bm{p}_{x}^{\text{ref}},\bm{p}_{y}^{\text{ref}},\bm{p}_{z}^{\text{ref}}]\in{\mathbb{R}}^{\text{16}}$, defines the reference contact state of each foot, $\bm{c}^{\text{ref}}\in{\mathbb{B}}^{\text{4}}$, and their reference Cartesian position in the world frame $x$-axis, $\bm{p}_{x}^{\text{ref}}\in{\mathbb{R}}^{\text{4}}$, $y$-axis, $\bm{p}_{y}^{\text{ref}}\in{\mathbb{R}}^{\text{4}}$, and $z$-axis, $\bm{p}_{z}^{\text{ref}}\in{\mathbb{R}}^{\text{4}}$ which are calculated online using the Raibert heuristic [48] to account for the current state of the robot. Throughout this paper, the limits enforced on the generation of $\bm{p}_{x}^{\text{ref}}$, $\bm{p}_{y}^{\text{ref}}$ and $\bm{p}_{z}^{\text{ref}}$ are $0.3$ m, $0.2$ m and $0.1$ m respectively from the nominal local foot position. An adjusted version of the BGS output, $\bm{\beta}_{G}$, is used for $\pi_{G}$ as not all the information in $\bm{\beta}_{L}$ is required. This has the form of $\bm{\beta}_{G}=[\bm{c}^{\text{ref}},\bm{p}_{z}^{\text{ref}},\Omega_{\text{stab}},\kappa]\in{\mathbb{R}}^{\text{10}}$ where $\Omega_{\text{stab}}\in\mathbb{R}$ characterises the inherent stability of a gait [5], and $\kappa\in\mathbb{B}$ is a logical flag to indicate a state of gait transition. We originally developed the BGS within [49] where the Froude number [5], $\Omega$, is used to trigger gait transition based exclusively on CoT which results in a set order of transitions. However, when applied to this work this method is not entirely suitable as now multiple biomechanics metrics and a set of auxiliary gaits need to be considered. One issue is that $\Omega$ values over $1$ are not compatible when calculating how many gait cycles, $C$, should a transition occur over. With this work investigating higher velocities than in [49], this has been resolved through calculating $C$ through
@@
This relationship ensures an almost instantaneous transition at a $\Omega$ of $\geq 2$, which is the typical value that quadruped animals transition to a run [5] instantaneously. Another limitation is that the calculation of the transition resolution, $n$, (how quickly a transition should be progressed each time step) only enables the transition between set gait pairs; this was not an issue in [49] as CoT efficiency was the only metric considered. As $\pi_{G}$ requires any gait transition pair to be possible, $\Omega_{\text{stab}}=g/hf^{2}$ [5] is utilised, where $g$ is gravity, $h$ is hip height and $f$ is gait frequency. Through the use of $\Omega_{\text{stab}}$, we are able to determine an indication of the inherent stability of any gait, hence a transition between a higher $\Omega_{\text{stab}}$ gait to a lower one should have smaller values of $n$ to increase the smoothness of the transition to promote stability. In the reverse scenario, a more harsh transition is more feasible hence larger values of $n$ should be produce for rapid transition. As such, $n$ is now calculated by
$n=1+\frac{\Omega_{\text{stab}}}{\Omega_{\text{stab}}^{\text{next}}}$ (3)

### v3

@@ I-C2 Actuator-structural Forces – Foot Contact Forces
Animals are observed to change gait to minimise actuator-structural forces (i.e. musculoskeletal forces) [37], which in biomechanics is measured through foot ground reaction force, $f_{\text{grf}}$. Similarly, $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{\tau_{\%}}$ reduce $f_{\text{grf}}$ through selecting the optimal gait for minimising $\tau_{\%}$. This supports that $\tau_{\%}$ is a suitable alternative to $f_{\text{grf}}$, which is further validated by a strong correlation between them within Figure 5b. However, not only does $\pi_{G}^{\tau_{\%}}$ maintain a trotting gait past optimal $f_{\text{grf}}$, but also the transition itself is instantaneous. In turn, $\pi_{G}^{\tau_{\%}}$ better reflects the animal data from [37] than $\pi_{G}^{\text{uni}}$. This could be explained by the metric being misinterpreted in [37]; with a high correlation between $\tau_{\%}$ and $f_{\text{grf}}$ with stability metric $c_{\text{avg}}^{\text{err}}$, it suggests instability causes transition, which is rare on flat terrain, as addressed in Section II.
@@ I-C3 Mechanical Work – External Work
Animals transition to preserve external mechanical work, $W_{\text{ext}}$, but do so before $\lambda^{W_{\text{ext}}}$, indicating relaxed minimisation (Figure 5a). $\pi_{G}^{\text{uni}}$ reflects this behaviour; transition occurs before $\lambda^{W_{\text{ext}}}$ yet minimal $W_{\text{ext}}$ is preserved. While $\pi_{G}^{W_{\text{ext}}}$ can reduce $W_{\text{ext}}$, its transition phase is extended over a larger range of velocities compared to animals, occurring just after $\lambda^{W_{\text{ext}}}$. In turn, this suggests that switching gaits between trotting and running offers minimal reductions in $W_{\text{ext}}$ resulting in a less definitive transition. However, $W_{\text{ext}}$ also seems to capture stability due to its high correlation with the stability metrics on $h_{\text{terr}}^{3}$ in Figure 5b, providing insight into its reduced role in Figure 5a where only flat terrain is present.
@@ I-C4 Stability – Stride Duration Coefficient of Variation
Within [35] animals are shown to reduce their stride duration coefficient of variation (CV) to preserve stability as high gait periodicity indicates stability. This behaviour is presented in Figure 5a, where animals are seen to initiate a transition phase when there is a considerable increase in stride CV, consequently leading to a decrease in CV and an increase in stability. $\pi_{G}^{\text{uni}}$ inherits the same strategy as only when a spike in stride CV is experienced does a transition phase begin resulting in improved stability through lowering stride CV. However, this is not the case with $\pi_{G}^{c^{\text{err}}}$ as it acts to reduce stride CV much more aggressively by mixing both slow, fast and auxiliary gaits which results in no clear transition phase being produced.
Only $\pi_{G}^{\text{uni}}$ consistently reflects all animal data sets and demonstrates the successful instillation of animal gait transition strategies. This also supports the notion that no singular metric can characterise animal gait selection and only through unification can similar behaviour in robots arise; the minimisation of the metrics is expected and is seen across all $\pi_{G}$ policies, but the intricacies of animal gait transition strategies are only seen in $\pi_{G}^{\text{uni}}$. Additionally, we have verified this behaviour transfers to real-world deployment within Supplementary Section 2; $\pi_{G}^{\text{uni}}$ achieves an $18\%$ and $30\%$ reduction in CoT compared to trotting and running on grassy terrain while also preventing the failure cases trotting exhibits.
Figure 6: Framework deployment in real-world environments to evaluate adaptability (a) The deployment of $\pi_{L}^{\text{bio}}$ on uneven grassy terrain with manual gait selection to cycle through all gaits with a $v_{\text{x}}^{\text{cmd}}$ of $0.5$ m/s where red shaded regions indicate a transition period. For subfigures (b), (c) and (d) our framework is deployed on loose timber, muddy grass and a low-friction board, with maximum velocity magnitudes of $1$ m/s, $2$ m/s and $1.7$ m/s respectively, where the event of critical loss of stability is indicated by the red dashed line, the bottom subplot uses the same gait colour code as Figure 4, and the heatmaps show the magnitudes of each metric increasing from purple to yellow. (e) Snapshots showing how animals also use a mixture of auxiliary gaits to overcome challenging terrain [48], where red circles are swing feet and green circles are stance feet.
@@ I-D Adaption to Real World Terrain
Although we have instilled animal gait transition strategies, gait procedural memory, and adaptive motion into our framework, its real-world proficiency without hardware deployment is uncertain. Grassy terrain may trap swing feet, and the ground often features irregularities. However, despite $\pi_{L}^{\text{bio}}$ only observing flat terrain during training, during hardware experiments it successfully realises all seven gaits on this terrain, as illustrated in Figure 6a, demonstrating that gait procedural memory and adaptive motion adjustment successfully transfer to real-world environments, providing a high level of adaptability, as shown in Supplementary Video S4.
Terrain that causes states of instability presents a substantial risk to the robot, hence we test the limits of our framework through deployment on loose timber, muddy grass, and a low-friction board, as presented in Figure 6b, Figure 6c and Figure 6d respectively and consolidated in Supplementary Video S5, with an additional experiment for external perturbations included within Supplementary Section 3. Each of the presented experiments showcases an off-nominal stability recovery event; however in the nominal scenario, the framework can maintain stability without changing gaits. In the case of loose timber, critical instability is caused when one rear foot slips on a plank, causing it to collide with another. In response, $\pi_{G}^{\text{uni}}$ utilises auxiliary gaits pronk and bound to recover, as depicted in Figure 6b. This strategy of utilising the auxiliary gaits for stability recovery is seen across all experiments and reflected in animals as highlighted in Figure 6e where a horse is observed to utilise bounding and limping gaits to traverse down complex rock formations.
In all experiments presented, a considerable increase in a combination of $W_{\text{ext}}$, $\tau_{\%}$ and $c_{\text{avg}}^{\text{err}}$ is experienced before a gait transition, while a weaker correlation is observed with CoT; spikes in $W_{\text{ext}}$, $\tau_{\%}$ and $c_{\text{avg}}^{\text{err}}$ directly coincide or even preempt gait changes while CoT peaks lag. This is expected with $c_{\text{avg}}^{\text{err}}$ and $W_{\text{ext}}$ as they capture periodicity and base height respectively. However, this is less expected for $\tau_{\%}$ to correlate with stability; this was never a factor investigated within [37], however, the correlation observed in these experiments and in Figure 5b provides strong evidence that this is the case.
@@ II Discussion
From taking inspiration from animal locomotion proficiency attributes, we have developed a locomotion framework capable of traversing complex and high-risk terrain despite the robot not utilising exteroceptive sensors nor experiencing any rough terrain during the training of $\pi_{L}^{\text{bio}}$. For $\pi_{L}^{\text{bio}}$, this is achieved through including the BGS output $\bm{\beta}_{L}$ (which encodes state dependent pseudo gait procedural memory and adaptive motion adjustments) within the observation space $\bm{o}_{L}$. This proves to be effective at overcoming rough terrain within Figure 3 as without the presence of $\bm{\beta}_{L}$ within $\bm{o}_{L}$, increased failure and instability are observed. This is the equivalent of removing an animal’s cerebellum functionality (resulting in reduced limb coordination and stability [45]) in turn supporting the claim that $\bm{\beta}_{L}$ effectively encodes pseudo gait procedural memory.
$\pi_{G}^{\text{uni}}$ for optimal gait selection greatly expands adaptability through instilling gait selection strategies used by animals. As demonstrated in Figure 6, $\pi_{G}^{\text{uni}}$ can maintain stability in the event of the terrain undergoing radical structural or friction coefficient adjustments. These scenarios pose risks to robots with vision systems, as they typically cannot detect ground friction or terrain changes beyond their front legs. Through the use of $\pi_{G}^{\text{uni}}$, this limitation is mitigated without implementing resource-heavy exteroceptive sensors. Comparing Figure 6b through 6d to Figure 6e showcases that animals and $\pi_{G}^{\text{uni}}$ utilise multiple auxiliary gaits to prevent failure. This behavior, untargeted during training, suggests that unifying these metrics encodes the intricacies of animal gait transitions.
One provoking observation is that actuator-structural forces appear to characterise instability. Considering that [37] validates by applying increased loads that could cause instability, it suggests this metric was initially misunderstood. Additionally, despite the employed biomechanics metrics only being tested on animals completing a linear forward trajectory on flat terrain, $\pi_{G}^{\text{uni}}$ upholds animal gait transition strategies across a wide range of terrains and base velocity commands. This supports that these metrics successfully characterise gait transitions and the notion that robots can indeed test biomechanics hypotheses, avoiding the resource, compatibility, and ethical challenges of animal testing. Moving forward, we aim to integrate our work with those that focus on Froude-free locomotion, such as [11, 12, 29] to achieve exemplary adaptability and efficiency at both Froude-characterised and Froude-free locomotion levels.
@@ III-A Control Framework Overview
At the core of this work, the Unitree A1 quadruped robot used in all experiments features $12$ degrees of freedom, $n$, which are all modeled as revolute joints, with their angular positions denoted as $\bm{q}\in{\mathbb{R}}^{n}$ and its base orientation represented as a rotation matrix $\bm{R}_{B}\in SO(\text{3})$. As discussed in Section I-A and outlined in Figure 1, both $\pi_{L}$ and $\pi_{G}$ are integrated within a control framework and supported by the SE and BGS for generation of the robot’s state data and gait references respectively. The final output of $\pi_{L}$ is target joint positions, $\bm{q}^{*}$, which are converted into joint torques, $\bm{\tau}^{*}$, through the following proportional derivative controller that get sent to the motors,
@@
where $K_{p}$ and $K_{p}$ are the proportional and derivative gains respectively. Throughout this work, a constant $K_{p}=25$ N/m and $K_{d}=1$ Ns/m are used while running at $1000$ Hz, while $\pi_{L}$ and $\pi_{G}$ are run at $500$ Hz and $100$ Hz respectively.
@@ III-B Bio-inspired Gait Scheduler
The BGS primary output, $\bm{\beta}_{L}=[\bm{c}^{\text{ref}},\bm{p}_{x}^{\text{ref}},\bm{p}_{y}^{\text{ref}},\bm{p}_{z}^{\text{ref}}]\in{\mathbb{R}}^{\text{16}}$, defines the reference contact state of each foot, $\bm{c}^{\text{ref}}\in{\mathbb{B}}^{\text{4}}$, and their reference Cartesian position in the world frame $x$-axis, $\bm{p}_{x}^{\text{ref}}\in{\mathbb{R}}^{\text{4}}$, $y$-axis, $\bm{p}_{y}^{\text{ref}}\in{\mathbb{R}}^{\text{4}}$, and $z$-axis, $\bm{p}_{z}^{\text{ref}}\in{\mathbb{R}}^{\text{4}}$ which are calculated online using the Raibert heuristic [49] to account for the current state of the robot. Throughout this paper, the limits enforced on the generation of $\bm{p}_{x}^{\text{ref}}$, $\bm{p}_{y}^{\text{ref}}$ and $\bm{p}_{z}^{\text{ref}}$ are $0.3$ m, $0.2$ m and $0.1$ m respectively from the nominal local foot position. An adjusted version of the BGS output, $\bm{\beta}_{G}$, is used for $\pi_{G}$ as not all the information in $\bm{\beta}_{L}$ is required. This has the form of $\bm{\beta}_{G}=[\bm{c}^{\text{ref}},\bm{p}_{z}^{\text{ref}},\Omega_{\text{stab}},\kappa]\in{\mathbb{R}}^{\text{10}}$ where $\Omega_{\text{stab}}\in\mathbb{R}$ characterises the inherent stability of a gait [5], and $\kappa\in\mathbb{B}$ is a logical flag to indicate a state of gait transition. We originally developed the BGS within [50] where the Froude number [5], $\Omega$, is used to trigger gait transition based exclusively on CoT which results in a set order of transitions. However, when applied to this work this method is not entirely suitable as now multiple biomechanics metrics and a set of auxiliary gaits need to be considered. One issue is that $\Omega>1$ values are not compatible when calculating how many gait cycles, $C$, should a transition occur over. With this work investigating higher velocities than in [50], this has been resolved through calculating $C$ through
@@
This relationship ensures an almost instantaneous transition at $\Omega\geq 2$, which is the typical value that quadruped animals transition to a run [5] instantaneously. Another limitation is that the calculation of the transition resolution, $\delta$, (how quickly a transition should be progressed each time step) only enables the transition between set gait pairs; this was not an issue in [50] as CoT efficiency was the only metric considered. As $\pi_{G}$ requires any gait transition pair to be possible, $\Omega_{\text{stab}}=g/hf^{2}$ [5] is utilised, where $g$ is gravitational field strength, $h$ is hip height and $f$ is gait frequency. Through the use of $\Omega_{\text{stab}}$, we are able to determine an indication of the inherent stability of any gait, hence a transition between a higher $\Omega_{\text{stab}}$ gait to a lower one should have smaller values of $\delta$ to increase the smoothness of the transition to promote stability. In the reverse scenario, a more harsh transition is more feasible hence larger values of $\delta$ should be produced for rapid transition. As such, $\delta$ is now calculated by
$\delta=1+\frac{\Omega_{\text{stab}}}{\Omega_{\text{stab}}^{\text{next}}}$ (3)